# Rate-Distortion Theory for Deductive Sources under Closure Fidelity

### 17 Apr 2026 v1

@@ Abstract
We study lossy compression of a finite statement source generated in a fixed deductive environment. The source symbols are statements in a knowledge base endowed with a proof system, and reconstruction fidelity is measured by preservation of deductive closure rather than by symbolwise equality. This induces, once the proof system and canonical order are fixed, a decomposition of the source into an irredundant core and redundant stored consequences. Under a natural disjointness condition on zero-distortion reconstruction sets, we show that the minimum zero-distortion rate equals the source mass of the core times the entropy of the source conditioned on that core. For reconstruction alphabets contained in the deductive closure of the source knowledge base, we further prove that the full rate–distortion function depends only on the core, so redundant states are invisible to both rate and distortion. When the decoder is limited to a bounded number of inference steps, we obtain an exact fixed-depth rate–delay–distortion characterization. Under an additional order-robustness assumption identifying the chosen core with the order-free essential set, this characterization interpolates between classical symbolwise compression and unconstrained deductive compression. These results formulate deductive compression as a structured source-coding problem and quantify how shared inference structure changes the fundamental limits of communication.
@@ I Introduction
The present formulation is related to, but distinct from, the broader semantic-communication and task-oriented communication literature, including early conceptual proposals, modern surveys, and recent mathematical formulations of semantic rate–distortion and semantic coding [6, 4, 19, 37, 31, 21, 39, 26, 12, 42, 22, 44, 28, 25, 46, 27, 15, 23, 33, 3]. Our formulation is narrower and more Shannon-theoretic: we fix a finite deductive environment and let the proof system itself induce zero-distortion equivalence through closure preservation.
@@
A further adjacent line comes from logic and database theory, especially the fixpoint semantics of Horn programs, finite-domain Datalog, and practical reasoning over compressed or materialized deductive stores [41, 7, 1, 11, 18, 13, 14]. We use this literature only as a proof-system substrate: the paper is not a database-theory paper, but an information-theoretic study of how shared inference changes source-coding limits.
@@
This paper asks three basic information-theoretic questions. First, what is the exact zero-distortion rate under closure fidelity? Second, how does the full rate–distortion function depend on the irredundant core? Third, how does this picture change when the decoder has only a limited inference budget? We answer these questions in a finite-alphabet setting. Our sharpest converse statements require an explicit disjointness assumption on zero-distortion reconstruction sets; we state this assumption up front because a central goal of the paper is to identify precisely where exact characterization is possible and where additional structure is still needed.
Our main message is that closure fidelity converts deductive redundancy into rate savings. In particular, the zero-distortion problem reduces to coding only the irredundant core, and the full rate–distortion function decomposes into a contribution from that core alone. We further show that when the decoder can perform at most $\delta$ inference steps, the relevant effective source becomes a $\delta$-irredundant core. Under an additional order-robustness assumption aligning the chosen core with the order-free essential set, this yields an exact rate–delay–distortion interpolation between classical symbolwise compression and unconstrained deductive compression. Source–channel separation, closure-adapted converse bounds, and heterogeneous-receiver extensions then follow as consequences of this source-model characterization.
@@
  • An exact rate–delay–distortion law for limited inference. When the decoder is allowed only a bounded inference budget, we introduce the $\delta$-irredundant core and prove the exact fixed-depth zero-distortion rate
@@
    together with the corresponding full rate–delay–distortion surface (Theorems 5.1 and 5.2). Under an additional order-robustness assumption ensuring $\text{Atom}(S_{O})=\operatorname{Ess}(S_{O})$, this yields a sharp interpolation between classical symbolwise coding and unconstrained deductive coding.
@@
The remainder of the paper is organized as follows. Section II introduces the deductive source model and closure fidelity. Section III proves the exact zero-distortion rate. Section IV derives the full rate–distortion decomposition over the irredundant core. Section V studies bounded inference and the resulting rate–delay–distortion tradeoff. Section VI presents consequences including source–channel separation, converse bounds, and a heterogeneous extension. Appendix A collects deferred proofs of the main results in Sections III–VI. Appendix B records concrete Datalog classes satisfying the standing assumptions. Appendix C contains additional examples and counterexamples.
@@ II-A Deductive Source and Irredundant Core
viewed as the standard logical closure operator associated with the proof system [40, 16, 17]. It satisfies:
@@ II-B Closure Fidelity and Closure Distortion
Section III studies the case $D=0$ and shows that the corresponding rate depends only on the irredundant core. Section V later replaces the unconstrained closure criterion by a bounded-inference version, yielding a rate–delay–distortion tradeoff.
@@ III Exact Zero-Distortion Rate
Using Theorem 3.1 and the classical identity $R(0;\,d_{H},\,P_{O})=H(P_{O})$ for zero Hamming distortion on a reconstruction alphabet containing $S_{O}$ [9, Sec. 10.5][10, Ch. 7], which is valid here because $S_{O}\subseteq\hat{S}_{O}$, completes the proof. ∎
@@
still follows from Proposition 3.1, but the converse must quantify the overlap structure among the sets $R_{\text{Cn}}(a)$, $a\in A$. This leads naturally to a confusability graph on the core. A graph-entropy-type characterization of that general case is, in our view, the key next theoretical breakthrough needed to complete the zero-distortion picture.
@@ IV Core Decomposition of the Full Rate–Distortion Function
V Limited Inference and the Rate–Delay–Distortion Tradeoff
Sections III and IV treat the decoder as having unrestricted deductive power. We now introduce a bounded inference budget. The resulting fidelity criterion distinguishes statements that are re-derivable eventually from those re-derivable within $\delta$ inference steps. This yields a natural rate–delay–distortion tradeoff.
@@ V-A Finite-Step Deduction and Derivation Depth
V-B Delay-Constrained Fidelity
Definition 5.3 ($\delta$-constrained closure distortion).
@@ V-B Delay-Constrained Fidelity
V-C Exact Rate–Delay–Distortion Laws
Definition 5.5 (Rate–delay–distortion function).
@@ V-C Exact Rate–Delay–Distortion Laws
Theorem 5.1 (Exact zero-distortion rate–delay function at fixed depth).
@@
Corollary 5.1 (Endpoints and monotonicity of the zero-distortion rate–delay function).
@@
Theorem 5.2 (Full rate–delay–distortion decomposition).
@@
The family $\{A_{\delta}\}_{\delta\geq 0}$ provides a quantitative depth-for-rate exchange. As $\delta$ increases, more states become reconstructible within the allowed inference budget and therefore disappear from the effective source. At $\delta=0$, nothing is free and the theory collapses to classical source coding. At $\delta\geq\mathsf{D_{d}}$, all eventually derivable shortcuts become free and, under the standing assumptions guaranteeing the large-$\delta$ endpoint identification, one recovers the deductive compression law of Section III.
@@ VI-A Separation and Operational Blocklength
Each distortion family considered here is a bounded single-letter distortion on a finite alphabet. Therefore the standard source–channel separation theorem for discrete memoryless sources and channels applies; see, e.g., [9, Chs. 7 and 10], [10, Ch. 3], [30, Chs. 7 and 8]. For $d_{\text{Cn}}$, the relevant source-coding function is $R_{\mathrm{sem}}(D;\,d_{\text{Cn}},\,P_{O})$; for $d_{\text{Cn}}^{\delta}$, it is $R_{\mathrm{sem}}(D,\delta;\,d_{\text{Cn}}^{\delta},\,P_{O})$. The zero-distortion thresholds follow by substituting the exact expressions from Theorems 3.1 and 5.1. Full details are given in Appendix A-F. ∎
Corollary 6.1 (Inference-budget thresholds from separation).
@@
    if $P_{A}H(\pi_{A})>\kappa C(W)$, then no finite inference budget suffices;
@@
Part (iv) follows because when equality never occurs, the two sets above coincide. Part (v) follows from $\phi(0)=H(P_{O})$. Part (vi) follows because $\phi(\delta)=P_{A}H(\pi_{A})$ for all $\delta\geq\mathsf{D_{d}}$, so if $P_{A}H(\pi_{A})>\kappa C(W)$, even the stabilized large-$\delta$ endpoint lies above channel capability. Part (vii) follows because $\phi(0)=H(P_{O})\geq\kappa C(W)$, while $\phi(\mathsf{D_{d}})=P_{A}H(\pi_{A})<\kappa C(W)$, and $\phi$ is non-increasing. ∎
@@
Part (i) is the standard Fano converse [9, Sec. 2.10][30, Sec. 1.9].
@@ VII Conclusion
This paper studied lossy compression of a finite knowledge source equipped with a fixed proof system, with fidelity measured by preservation of deductive closure rather than by symbolwise equality. The central conclusion is that, under closure fidelity, the effective source complexity is governed by the irredundant core of the knowledge base rather than by the full stored source. Under an explicit disjointness condition on zero-distortion reconstruction sets, we showed that the zero-distortion rate is exactly $P_{A}H(\pi_{A})$, and that the full rate–distortion function decomposes into a contribution from that core alone whenever the reconstruction alphabet is contained in $\text{Cn}(S_{O})$. We further derived an exact fixed-depth rate–delay–distortion characterization for decoders with limited inference budgets; under an additional order-robustness assumption ensuring that the chosen core coincides with the essential set, this yields a sharp interpolation between classical symbolwise compression and unconstrained deductive compression.
@@ Appendix A Proofs of Sections III–VI
Let $P_{a}:=P_{\hat{S}\mid S}(\cdot\mid a)$ for $a\in A$, let
$\bar{P}_{A}:=\sum_{a\in A}\pi_{A}(a)\,P_{a},$
and let $Q:=P_{\hat{S}\mid T}(\cdot\mid*)$ denote the output law assigned to the redundant class. Then the output marginal is
@@
$\geq P_{A}\Bigl(H(\bar{P}_{A})-\sum_{a\in A}\pi_{A}(a)H(P_{a})\Bigr).$
The bracketed term is exactly $I(\mathsf{A}_{o};\hat{\mathsf{S}}_{o})$ for the core sub-source $\mathsf{A}_{o}\sim\pi_{A}$ and test channel $P_{\hat{S}\mid A}$. Hence every feasible global test channel satisfies
@@
for all $(s_{o},\hat{s}_{o})$, because allowing more inference steps can only enlarge the zero-distortion feasible set. Therefore the feasible test-channel set for zero $\delta_{2}$-distortion contains that for zero $\delta_{1}$-distortion, and thus
@@
Let $p_{A}:=\epsilon_{\text{Cn}}/P_{A}$. Applying the classical Fano inequality [9, Sec. 2.10][30, Sec. 1.9] to the $|A|$-ary source conditioned on $B=1$ gives
@@ C-C A Minimal Exact Zero-Distortion Example
C-D A Simple Rate–Delay Example
@@
Then $d$ is derivable from $a$ in two proof steps, and $e$ is derivable from $b$ in two proof steps. The irredundant core is
@@
However, with only one proof step available, neither $d$ nor $e$ is redundant relative to $S_{O}\setminus\{d\}$ or $S_{O}\setminus\{e\}$, because the intermediate statements $c$ and $f$ are not stored. Hence

### 20 Apr 2026 v2

@@ Abstract
We study lossy compression of a finite statement source generated in a fixed deductive environment. The source symbols are statements in a knowledge base endowed with a proof system, and reconstruction fidelity is measured by preservation of deductive closure rather than by symbolwise equality. This induces, once the proof system and canonical order are fixed, a decomposition of the source into an irredundant core and redundant stored consequences. Under a natural disjointness condition on zero-distortion reconstruction sets, we show that the minimum zero-distortion rate equals the source mass of the core times the entropy of the source conditioned on that core. For reconstruction alphabets contained in the deductive closure of the source knowledge base, we further prove that the full rate–distortion function depends only on the core, so redundant states are invisible to both rate and distortion. When the decoder is limited to a bounded inference-depth budget, we obtain an exact rate–depth–distortion characterization. Under an additional order-robustness assumption identifying the chosen core with the order-free essential set, this characterization interpolates between classical symbolwise compression and unconstrained deductive compression. These results formulate deductive compression as a structured source-coding problem and quantify how shared inference structure changes the fundamental limits of communication.
@@ I Introduction
The present formulation is related to, but distinct from, the broader semantic-communication and task-oriented communication literature, including early conceptual proposals[4], modern surveys[10], and recent mathematical formulations of semantic rate–distortion and semantic coding [18, 20, 32, 13, 19]. Our formulation is narrower and more Shannon-theoretic: we fix a finite deductive environment and let the proof system itself induce zero-distortion equivalence through closure preservation.
@@
A further adjacent line comes from logic and database theory, especially the fixpoint semantics of Horn programs, finite-domain Datalog, reasoning over materialized deductive stores, and the broader theory of incomplete information and certain answers [29, 5, 1, 9, 15, 12, 11]. Our redundant/materialized split is closest to the materialized-view perspective: the set $J$ behaves as stored consequences that need not be communicated once the irredundant premises are available. An incomplete-information analogue would replace deductive closure by a certain-answer operator, leading to a semantics-based fidelity notion. We do not pursue that extension here; the present paper remains a Shannon-theoretic study of a fixed deductive environment.
@@
This paper asks three basic information-theoretic questions. First, what is the exact zero-distortion rate under closure fidelity? Second, how does the full rate–distortion function depend on the irredundant core? Third, how does this picture change when the decoder has only a limited inference-depth budget (a limited number of iterations)? We answer these questions in a finite-alphabet setting. Our sharpest converse statements require an explicit disjointness assumption on zero-distortion reconstruction sets; we state this assumption up front because a central goal of the paper is to identify precisely where exact characterization is possible and where additional structure is still needed.
Our main message is that closure fidelity converts deductive redundancy into rate savings. In particular, the zero-distortion problem reduces to coding only the irredundant core, and the full rate–distortion function decomposes into a contribution from that core alone. We further show that when the decoder can perform at most $\delta$ iterations of the immediate-consequence operator, the relevant effective source becomes a $\delta$-irredundant core. Under an additional order-robustness assumption aligning the chosen core with the order-free essential set, this yields an exact rate–depth–distortion interpolation between classical symbolwise compression and unconstrained deductive compression. Source–channel separation, closure-adapted converse bounds, and heterogeneous-receiver extensions then follow as consequences of this source-model characterization.
@@
  • An exact rate–depth–distortion law for limited inference. When the decoder is allowed only a bounded inference-depth budget, we introduce the $\delta$-irredundant core and prove the exact zero-distortion rate–depth law
@@
    together with the corresponding full rate–depth–distortion function $R_{\mathrm{sem}}(D,\delta)$ over $(D,\delta)$ (Theorems 5.1 and 5.2). Under an additional order-robustness assumption ensuring $\text{Atom}(S_{O})=\operatorname{Ess}(S_{O})$, this yields a sharp interpolation between classical symbolwise coding and unconstrained deductive coding.
@@
The remainder of the paper is organized as follows. Section II introduces the deductive source model and closure fidelity. Section III proves the exact zero-distortion rate. Section IV derives the full rate–distortion decomposition over the irredundant core. Section V studies bounded inference and the resulting rate–depth–distortion tradeoff. Section VI presents consequences including source–channel separation, converse bounds, and a heterogeneous extension. Appendix A collects deferred proofs of the main results in Sections III–VI. Appendix B records concrete Datalog classes satisfying the standing assumptions. Appendix C contains additional examples and counterexamples.
@@ II-A Deductive Source and Irredundant Core
viewed as the standard logical closure operator associated with the proof system [14]. It satisfies:
@@ II-B Closure Fidelity and Closure Distortion
Section III studies the case $D=0$ and shows that the corresponding rate depends only on the irredundant core. Section V later replaces the unconstrained closure criterion by a bounded-inference version, yielding a rate–depth–distortion tradeoff.
@@ III Exact Zero-Distortion Rate
Using Theorem 3.1 and the classical identity $R(0;\,d_{H},\,P_{O})=H(P_{O})$ for zero Hamming distortion on a reconstruction alphabet containing $S_{O}$ [7, Sec. 10.5][8, Ch. 7], which is valid here because $S_{O}\subseteq\hat{S}_{O}$, completes the proof. ∎
@@
still follows from Proposition 3.1, but the converse must quantify the overlap structure among the sets $R_{\text{Cn}}(a)$, $a\in A$. This leads naturally to a confusability graph on the core. A graph-entropy-type characterization of that general case is, in our view, the key next theoretical breakthrough needed to complete the zero-distortion picture. A related analysis for ambiguous alphabets under asymptotic equipartition is given in [28].
@@ IV Core Decomposition of the Full Rate–Distortion Function
V Limited Inference and the Rate–Depth–Distortion Tradeoff
Sections III and IV treat the decoder as having unrestricted deductive power. We now introduce a bounded inference-depth budget. The resulting fidelity criterion distinguishes statements that are re-derivable eventually from those re-derivable within $\delta$ iterations of $T_{\mathsf{PS}}$. This yields a natural rate–depth–distortion tradeoff.
@@ V-A Finite-Step Deduction and Derivation Depth
V-B Depth-Constrained Fidelity
Definition 5.3 ($\delta$-step closure distortion).
@@ V-B Depth-Constrained Fidelity
V-C Exact Rate–Depth–Distortion Laws
Definition 5.5 (Rate–depth–distortion function).
@@ V-C Exact Rate–Depth–Distortion Laws
Theorem 5.1 (Exact zero-distortion rate–depth tradeoff).
@@
Corollary 5.1 (Endpoints and monotonicity of the zero-distortion rate–depth function).
@@
Theorem 5.2 (Full rate–depth–distortion decomposition).
@@
The family $\{A_{\delta}\}_{\delta\geq 0}$ provides a quantitative depth-for-rate exchange. As $\delta$ increases, more states become reconstructible within the allowed depth budget and therefore disappear from the effective source. At $\delta=0$, nothing is free and the theory collapses to classical source coding. At $\delta\geq\mathsf{D_{d}}$, all eventually derivable shortcuts become free and, under the standing assumptions guaranteeing the large-$\delta$ endpoint identification, one recovers the deductive compression law of Section III.
@@ VI-A Separation and Operational Blocklength
Each distortion family considered here is a bounded single-letter distortion on a finite alphabet. Therefore the standard source–channel separation theorem for discrete memoryless sources and channels applies; see, e.g., [7, Chs. 7 and 10], [8, Ch. 3], [22, Chs. 7 and 8]. For $d_{\text{Cn}}$, the relevant source-coding function is $R_{\mathrm{sem}}(D;\,d_{\text{Cn}},\,P_{O})$; for $d_{\text{Cn}}^{\delta}$, it is $R_{\mathrm{sem}}(D,\delta;\,d_{\text{Cn}}^{\delta},\,P_{O})$. The zero-distortion thresholds follow by substituting the exact expressions from Theorems 3.1 and 5.1. Full details are given in Appendix A-F. ∎
Corollary 6.1 (Depth-budget thresholds from separation).
@@
    if $P_{A}H(\pi_{A})>\kappa C(W)$, then no finite depth budget suffices;
@@
Part (iv) follows because when equality never occurs, the two sets above coincide. Part (v) follows from $\phi(0)=H(P_{O})$. Part (vi) follows because $\phi(\delta)=P_{A}H(\pi_{A})$ for all $\delta\geq\mathsf{D_{d}}$, so if $P_{A}H(\pi_{A})>\kappa C(W)$, even the stabilized large-$\delta$ endpoint lies above channel capability. Part (vii) follows because $\phi(0)=H(P_{O})>\kappa C(W)$, so neither threshold is attained at $\delta=0$, whereas $\phi(\mathsf{D_{d}})=P_{A}H(\pi_{A})<\kappa C(W)$; monotonicity of $\phi$ then gives
@@
Part (i) is the standard Fano converse [7, Sec. 2.10][22, Sec. 1.9].
@@ VI-C One Heterogeneous Extension
If $P_{A}=0$, the claim is immediate by constant reconstruction to any representative of a core symbol guaranteed by (H1). Assume henceforth that $P_{A}>0$.
@@ VII Conclusion
This paper studied lossy compression of a finite knowledge source equipped with a fixed proof system, with fidelity measured by preservation of deductive closure rather than by symbolwise equality. The central conclusion is that, under closure fidelity, the effective source complexity is governed by the irredundant core of the knowledge base rather than by the full stored source. Under an explicit disjointness condition on zero-distortion reconstruction sets, we showed that the zero-distortion rate is exactly $P_{A}H(\pi_{A})$, and that the full rate–distortion function decomposes into a contribution from that core alone whenever the reconstruction alphabet is contained in $\text{Cn}(S_{O})$. We further derived an exact rate–depth–distortion characterization under a bounded inference-depth budget; under an additional order-robustness assumption ensuring that the chosen core coincides with the essential set, this yields a sharp interpolation between classical symbolwise compression and unconstrained deductive compression.
@@ Appendix A Proofs of Sections III–VI
Let $P_{a}:=P_{\hat{S}\mid S}(\cdot\mid a)$ for $a\in A$, and let
$\bar{P}_{A}:=\sum_{a\in A}\pi_{A}(a)\,P_{a}.$
The quantity
$H(\bar{P}_{A})-\sum_{a\in A}\pi_{A}(a)H(P_{a})$
is exactly $I(\mathsf{A}_{o};\hat{\mathsf{S}}_{o})$ for the core sub-source $\mathsf{A}_{o}\sim\pi_{A}$ and test channel $P_{\hat{S}\mid A}$.
If $0<P_{A}<1$, let
$Q:=\sum_{j\in J}\frac{P_{O}(j)}{P_{J}}\,P_{\hat{S}\mid S}(\cdot\mid j).$
Then the output marginal is
@@
$\geq P_{A}\Bigl(H(\bar{P}_{A})-\sum_{a\in A}\pi_{A}(a)H(P_{a})\Bigr)$
$=P_{A}\,I(\mathsf{A}_{o};\hat{\mathsf{S}}_{o}).$
If $P_{A}=1$, then $J=\varnothing$ and $T=S_{o}$ almost surely. The output marginal is simply $\bar{P}_{A}$, so
$I(T;\hat{S}_{o})$ $=H(\bar{P}_{A})-\sum_{a\in A}\pi_{A}(a)H(P_{a})$
$=I(\mathsf{A}_{o};\hat{\mathsf{S}}_{o})=P_{A}\,I(\mathsf{A}_{o};\hat{\mathsf{S}}_{o}).$
Therefore every feasible global test channel satisfies
@@
for all $(s_{o},\hat{s}_{o})$, because allowing more iterations (of $T_{\mathsf{PS}}$) can only enlarge the zero-distortion feasible set. Therefore the feasible test-channel set for zero $\delta_{2}$-distortion contains that for zero $\delta_{1}$-distortion, and thus
@@
Let $p_{A}:=\epsilon_{\text{Cn}}/P_{A}$. Applying the classical Fano inequality [7, Sec. 2.10][22, Sec. 1.9] to the $|A|$-ary source conditioned on $B=1$ gives
@@ C-C A Minimal Exact Zero-Distortion Example
C-D A Simple Rate–Depth Example
@@
Then $d$ is derivable from $a$ with derivation depth $2$, and $e$ is derivable from $b$ with derivation depth $2$. The irredundant core is
@@
However, with depth budget $\delta=1$, neither $d$ nor $e$ is redundant relative to $S_{O}\setminus\{d\}$ or $S_{O}\setminus\{e\}$, because the intermediate statements $c$ and $f$ are not stored. Hence